if 32.0 g CaC2 are consumed in this reaction, how many moles of H20 are needed?

1
Expert's answer
2022-03-25T05:57:06-0400

The reaction equation is:

CaC₂ + 2H₂O → Ca(OH)₂ + C₂H₂

Moles of CaC₂ = 32 / (40 + 12 x 2)

Moles of CaC₂ = 0.5

Molar ratio of CaC₂ to H₂O is 1:2

Moles of H2O = 0.5×2 = 1.0